OpenStudy (anonymous):

For a satellite at a height of 22,300 miles, at what speed must the satellite be traveling to achieve a circular orbit?

OpenStudy (irishboy123):

balancing forces \(F = \dfrac{GMm}{r^2} = \dfrac{mv^2}{r}\) \(v = \sqrt{\dfrac{GM}{r}}\) \( \mu = GM = ( 398\ 600.4418 \pm 0.0008 ) {\rm \ km^{3} \ s^{-2} } \)  r is the distance between centres of mass...

OpenStudy (irishboy123):

that's doesn't answer your question but its gives you what you need to think about